### v2

@@ 5 Conclusion
In this paper, we have argued that existing tests for choice overload are underpowered, and that this may explain the confusing and contradictory picture that emerges from the current literature. We proposed that, by collecting more data and making use of theory, one can design better tests. We have demonstrated that, in a novel data set, our approach finds choice overload. Indeed, our new data speak so loudly that with hindsight, our methodological innovations would not have been necessary to detect choice overload in our own data set. We believe that the innovations are of interest nonetheless, and we also note that such things are bound to occur if –as we did– one genuinely designs the empirical strategy before going into the field.
One unexpected finding from our experiment is that choice appears to violate the RUM even in smaller choice sets. This is perhaps surprising given the relative simplicity of the choice items. Understanding the cause of these violations strikes us as an interesting avenue for future research. For our purposes, it suggests that RUM may not be an appropriate benchmark against which to look for choice overload, at least in this experimental task.
@@ Appendix A Supplementary Appendix: Experimental Details
Here we present the experiment instructions, quiz, and details about the choice objects.
@@ A.1 Instructions and Quiz
Subjects were first shown this instructions screen before proceeding to the quiz (below). Subjects were given two attempts at the quiz.
[Uncaptioned image] [Uncaptioned image]
@@ A.2 Choice Alternatives with Summary Data
Table 1 gives full list of the choice objects with their values in experimental points.
Table 1: List of Options with Values
@@
Table 2 shows how often each choice set was shown to subjects and how often the default was chosen from it. Choice sets are given by the alternatives they contain: $[0,1,2]$ represents the choice set with the default (0) and options with IDs 1 and 2 from Table 1.
@@
Appendix B Implementing recent computational innovations
We implemented the computational procedure in Smeulders et al. (2020). To our knowledge, this is the first such implementation beyond their own illustrative example. The implementation and some not entirely obvious modifications are described next.
The procedure is motivated by the fact that computation of the matrix $\bm{A}$ and also computation (3.4) is hard, and yet the latter needs to be repeated many times. It exploits that, because $\bm{BA}$ has many more columns than rows, there always exists a sparse (in the loose sense of having relatively few nonzero entries) $\arg\max$ to problem (3.4). We will avoid solving (3.4) as stated, or ever computing $\bm{BA}$ (though the latter is feasible here), by guessing the nonzero entries. Formally, this goes as follows. (We drop $N$ for ease of notation.):

### v3

@@ 5 Conclusion
In this paper, we argued that existing tests for choice overload are not very sensitive, potentially explaining the ambiguous picture that emerges from the current literature. We proposed that, by collecting more data and fully using restrictions from economic theory, one can design better tests. We find choice overload in a novel data set, while standard approaches would only have had a 19% chance of doing so. Indeed, our new data speak so loudly that with hindsight, our econometric innovations would not have been necessary to detect choice overload. We believe that the innovations are of interest nonetheless, and we also note that such things are bound to occur if one genuinely designs the empirical strategy before collecting data (as we did).
@@ A.1 Instructions and Quiz
Subjects were first shown an instructions screen before proceeding to the quiz; see Figure 3. Subjects were given two attempts.
Figure 3: Instructions.
Figure 4: Quiz screen 1.
Figure 5: Quiz screen 2.
Figure 6: Quiz screen 3.
@@ A.2 Choice Alternatives with Summary Data
Table 1 gives a full list of the choice objects with their values in experimental points. Table 2 shows how often each choice set was shown to subjects and how often the default was chosen from it. Choice sets are given by the alternatives they contain: e.g. $[0,1,2]$ represents the choice set with the default (0) and options with IDs 1 and 2 from Table 1. Figure 7 shows a histogram of the default choice probabilities of all choice sets.
@@
Table 1: List of Options with Values
@@
Appendix B Supplementary Appendix:
Implementing Recent Computational Innovations
In the process of writing this paper, we implemented the computational procedure in Smeulders et al. (2021). To our knowledge, this is the first such implementation beyond their own illustrative example. The implementation and some not entirely obvious modifications are described next. This material is not in the main text because we did not end up using the implementation in the empirical work.
Smeulders et al.’s (2021) procedure is motivated by the fact that computation of the matrix $\bm{A}$ and also computation (3.4) is hard, and yet the latter needs to be repeated many times. It exploits that, because $\bm{BA}$ has many more columns than rows, there always exists a sparse (in the loose sense of having relatively few nonzero entries) $\arg\max$ to problem (3.4). We will avoid solving (3.4) as stated, or ever computing $\bm{BA}$ (though the latter is feasible here), by guessing the nonzero entries. Formally, this goes as follows. (We drop $N$ for ease of notation.):